Reports 1-1 of 1 Clear search Modify search
AdV-INF (Environmental noise mitigation)
fiori, tringali, passaquieti - 12:04 Wednesday 15 December 2021 (54224) Print this report
Acoustic injections NEB HVAC

During yesterday maintenance we performed some acoustic injections in the NEB AHU room using one subwoofer loudspeaker.

At 9:11 UTC we switched off the AHU. The reference quiet time is between 9:12:30 and 9:17:30 UTC. Then we injected a 30s cycling sweep noise with frequency increasing linearly from 10 Hz to 100Hz. The good time of sweeps is between 9:19 and 9:22 UTC. After that euroterm people entered the AHU room and we stopped measuring.

Figure 1 is a spectrogram of the microphone inside the AHU room and the NOISE channel used for the injection.

Figure 2 shows the noise increase in some selected sensors. Seems that the acoustic noise in the hall reacted sensibly to this noise injection, in fact the excess acoustic noise in the hall is more or less of the same magnitude as that injected in the AHU room. A similar excess noise is seen by the SAS microphone and accelerometers on the ducts. It is possible that our injection excited different paths. We have to study better the data, also checking for coherences. We also measure a noticeable increase of AHU floor vibrations: this could be vibrations of the loudspeaker box. Next time we have to better seismically isolate the speaker.

We plan to complete our acoustic injection plan in another dedicated shift.

Images attached to this report
Comments to this report:
fiori, tringali, passaquieti, paoletti - 17:34 Tuesday 21 December 2021 (54308) Print this report

We performed some more sweep and white band limited noise in NEB:

December 17:

we seismically isolate the loudspeaker and repeated the sweep injection (AHU off)

Today (December 21):

we repeated the same injections with the (seismically isolated) loudspeakers inside the AHU fan box (AHU off)

The attached log file lists the single actions and gps.

The attached Figure 1 shows the effect of the sweep injection of today on some relevant environmental sensors.

The attached Figure 2 compares the effect of white noise acoustic injection with the loudspeaker into the AHU room (golden curve) and the loudspeaker inside the AHU fan box (blue) (purple is quiet time with the AHU off).


Images attached to this comment
Non-image files attached to this comment
fiori, tringali, paoletti - 17:41 Tuesday 21 June 2022 (56216) Print this report

Today we reinstalled the acoustic setup at the NEB AHU room. We are ready to perform injections when we get green light from the commissioning. The purpose is to remeasure acosutci couplings after the acoustic isolation works of the AHU room and the ducts.

Search Help